Proposal:
K Kids
Target of Oportunity:
No
Time Observed:
11.5
Time Engineering:
0.0
Time lost technical:
0.5
Time lost weather:
0.0
Time lost other:
0.0
Primary Investigator:
Todd Henry
Institution:
SMARTS
Instrument:
CTIO 1.5m Echelle
Organization:
SMARTS
Total Program Time:
12

ThAr lamps looks like BIAS e.g., 1119, 21, 22 etcetera..
Marco Bonati tested and found no apparent failure, so, resumed normal operation.
File 1129 T0349576261 acquired in wrong mode (while testing lamps, shutter, system.!)
Underline: If we 'only' use SEND from NOS, lamps fails (standard procedure fails.) In the tests I did, it also failed if sent from "preset modes / scripts".
Found this workaround: ON the lamp (@ PI_lamps.vi) + SEND @ NOS: worked!
Observation: Once lamp has been taken, PI_LAMPS shows the system turns it OFF. Ok.
File 1185 (GJ1224): telescope moved SW. Sent 2 new integrations. Is not windy but telescope is pointing N, against the wind.
No time for: T0354944123, T0197959526, HD223807.
Last integration on 'HIP111909' ended with the Sun at -9.5 degrees of altitude.